find slope of -4x +5y =10 and slope of 3x + 4y = 0
are they perpendicular, parallel or neither

Answer by Alan3354(69443)

find slope of -4x +5y =10 and slope of 3x + 4y = 0
are they perpendicular, parallel or neither
-------------
-4x +5y =10
--
Solve for y.
You get y = mx + b
The slope is the coefficient of x, m = slope.
-----
If the 2 slopes are the same, they're parallel.
If m1 = -1/m2 they're perpendicular.